OpenStudy (anonymous):

Solve y=2x-1 2x-y-1=0

OpenStudy (anonymous):

Replace (2x-1) in the place of y in second equation.

OpenStudy (anonymous):

they ask if the system has exactly one solution or if it has infinitely many solutions. which would be my answer for this equation?

OpenStudy (anonymous):

Infinitely many solutions

OpenStudy (anonymous):

show the steps please

OpenStudy (anonymous):

y=2x-1 2x-y-1=0 2x-(2x-1)-1=0 2x-2x+1-1=0 2x-2x=0 So infinitely many solutions

OpenStudy (anonymous):

2x-2x=0 2x=2x

OpenStudy (anonymous):

why does the 2x=2x if right before that it was 2x-2x=0? wouldn't it be 0=0?

OpenStudy (anonymous):

2x-2x=0 0=0 If we add both sides by 2x then.. 2x-2x+2x=0+2x 2x=2x

OpenStudy (anonymous):

why did you need to add the] 2x?

OpenStudy (anonymous):

It's not needed, 2x-2x=0 is enough

OpenStudy (anonymous):

If it wasn't needed, why'd you add 2x to -2x and to the 0 ?

OpenStudy (anonymous):

If we add 2x also it doesn't matter. It was just another way of showing that they can have infinite values.

OpenStudy (anonymous):

How did you know that it was gonna be infinite value?

OpenStudy (anonymous):

2x=2x give any value to x, it's same.

OpenStudy (anonymous):

in the beginning why did we have to Replace (2x-1) in the place of y in second equation. ?

OpenStudy (anonymous):

i need step by step instructions.

OpenStudy (anonymous):

If there are two variables with one equation it's not possible to get the answer. But in this question there are two equation's only using the variables x and y. So first try to write x in terms of y or y in terms of x. For Eg. 4x-y=4 You cannot solve the above equation without knowing x or y value or x in terms of y or vice versa. For above equation if it's given that y=2x then replace 2x in place of y so there is only on variable, and you can solve. 4x-2x=4 2x=4 x=2 y=2x=4 Understood?

OpenStudy (anonymous):

Ohkay I get it with that example. show me another?

OpenStudy (anonymous):

You try to make a question and solve..

OpenStudy (anonymous):

Why me? I'm the one that's lost/:

OpenStudy (anonymous):

It would improve your ability to make question's. You try to make a question similar to the one you posted.

OpenStudy (anonymous):

3x+y=4 -3x=y-7

OpenStudy (anonymous):

what would be step by step for this one?

OpenStudy (anonymous):

You try to do, I will say, if you made any mistake.

OpenStudy (anonymous):

First try to make x in terms of y or vice versa

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

3x+y=4 y=4-3x Substitute this value in second equation

OpenStudy (anonymous):

I'm so dumb -_- I don't get what you did to the second equation. I mean like i just dont know how you do that to ever probelm..

OpenStudy (anonymous):

yvette, what you need to do is notice that the second equation is all setup with a y=... so you can use what y equals to substitute into the first equation. fror example with the one you just posted: 3x+y=4 y=4-3x you can substitute (4-3x) for the y in the first equation... so... 3x + (4-3x) = 4 then continue to solve... 3x + 4 - 3x = 4 then you should notice that both sides can be made to be the same: 3x + 4 = 3x + 4